The equation of state for 5 g of oxygen at a pressure P and temperature T, when occupying a volume V, will be


        (a)    PV = (5/16)RT   


(b)    PV = (5/32) RT


        (c)    PV = 5 RT


        (d)    PV = (5/2) RT


where R is the gas constant.


Grade:10

1 Answers

Amit Saxena
10 years ago

(b)
n = 5/32
:. PV = 5/32 RT [? PV = nRT]